OpenStudy (anonymous):

solve 2log(x+11)=(1/2)^x.

OpenStudy (anonymous):

what's log's base?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Oh okay...

OpenStudy (anonymous):

\[2\log (x+ 11) = \left( \frac{1}{2}\right)^x\]

OpenStudy (anonymous):

that is the question right?

OpenStudy (anonymous):

Yeah, let me show you what I've got so far.

OpenStudy (anonymous):